OpenStudy (anonymous):

Solve. 7^(x – 5) = 7^(4x + 7)

OpenStudy (jdoe0001):

$$ 7^{x – 5} = 7^{4x + 7} \ ? $$

OpenStudy (anonymous):

yes!But I think I figured it out. The bases are already identical so i can eliminate them. Then i just solve x-5 = 4x+7 and the final answer is x=-4

OpenStudy (jdoe0001):

yes :)

OpenStudy (anonymous):

:) thanks!